# Cache Coherence Over Disaggregated Memory

### v3

@@ 6.3. Index Support
With the abstraction layer enabled by SELCC, we can easily transition existing database indexes into disaggregated indexes with the following two steps. The first step is organizing the basic data structure into Global Cache Lines (GCL). For data structures that already organize multiple data objects into blocks, e.g., B-tree, R-tree, ART tree, and hash table, this process is simplified by aligning the node structures onto GCL. In cases where the original data structure is in-memory and does not organize multiple data objects into blocks (e.g., skip list), there are two options: Reorganizing multiple data objects into blocks based on their locality, or adjusting the global cache line size to match a single data object. The former approach minimizes space overhead, while the latter requires no code modifications. The second step involves replacing local shared-exclusive latches in the algorithm with SELCC_XLock/SELCC_SLock. SELCC ensures both read-write atomicity and cache coherence globally.
We implement a concurrent B-link-tree over SELCC with around 1,000 lines of code and evaluate its performance using the YCSB benchmark (Fig. 16). It achieves competitive performance compared to state-of-the-art B-trees over shared disaggregated memory.
@@ 6.4. Transaction Support
We can migrate existing concurrency control algorithms to the disaggregated architecture by leveraging the abstraction layer enabled by SELCC as follows. (1) Tuples should be properly organized into GCLs. (2) Local shared-exclusive latches are replaced with SELCC_XLock/SELCC_SLock locks. (3) Algorithms that require monolithic timestamps utilize Atomic provided by SELCC API to perform RDMA Fetch-and-Add (FAA) operations on a global timestamp generator to obtain monotonically increasing timestamps.
We implement three types of algorithms over SELCC: Two-Phase Locking with no wait strategy (2PL), Timestamp Ordering (TO), and Optimistic Concurrency Control (OCC). Tuples are organized in a heap style, meaning they are placed in GCLs in chronological order of insertion. To ensure atomicity of tuple accesses, these accesses must be protected by SELCC_XLock/ SELCC_SLock locks. For two-phase locking, SELCC latches on the GCLs are reused for locking purposes, minimizing the RDMA round trips required by the transaction concurrency control.
To support durability in disaggregated memory, durable storage media is leveraged for write-ahead logging. Since all transactions are executed within the same compute node via RDMA, transaction support over SELCC does not require two-phase commit protocols.
@@ 7.1. Evaluating SELCC
Baselines. To show the efficiency of SELCC, we compare SELCC against three competitors over disaggregated memory: (1) GAM (Cai et al., 2018), an RPC-based cache-coherence protocol designed for distributed shared memory. We test GAM with different consistency models: total store order consistency and sequential consistency, corresponding to GAM (TSO) and GAM (SEQ) in the figures. (2) ScaleStore (Ziegler et al., 2022), state-of-the-art RPC-based protocol designed for distributed shared memory over distributed shared SSDs. To adapt it for disaggregated memory, we configure it differently for compute and memory nodes. For the memory nodes, the buffer pool sizes are sufficiently large to fully store the data in memory, eliminating SSD access [^4] There is one page provider and one message handler on each memory node by default.. For the compute nodes, the system no longer allocates data and instead functions as a stateless compute server. The last baseline is SEL (Ziegler et al., 2023b), a one-sided access framework that operates without compute-side caching. While it employs the shared-exclusive latch (SEL) to ensure RDMA access atomicity, it bypasses the cache coherence problem by disabling caching, which can lead to low performance.
@@
Benchmarks. We test the competitors by the micro-benchmark tool that allows for adjustments in sharing ratios, read/write ratios, data skewness and access locality. Each compute server issues 16 Million accesses over 24 Million allocated GCLs (2KB per GCL, 48GB in total) The overall throughput with different read ratios 100% (Read only), 95% (Read intensive), 50% (Write intensive), 0% (Write only) are tested. Read only workload has similar result to read intensive workload, so we omit its result in the presentation.
@@ 7.1.1. Evaluating the Scalability of SELCC
To evaluate the scalability of the SELCC protocol, we run the benchmark under a uniformly distributed workload while varying the number of compute nodes. To avoid bottlenecks due to memory nodes’ RDMA bandwidth, we scale the number of memory nodes in proportion to the number of compute nodes. Each compute node executes 8 local threads over 16GB local cache. We compare SELCC under various sharing ratios ($sr$), following the methodology in (Yang et al., 2024; Cai et al., 2018; Depoutovitch et al., 2023). The sharing ratio ($sr$) indicates the percentage of allocated data accessible by all compute nodes, while the remainder is accessed privately. When the sharing ratio is zero, the system essentially operates as a sharding-based system over disaggregated memory.
Fig. 12 shows the experimental results. The point values represent the overall throughput while the bar values indicate the proportion of operations requiring invalidation messages. For the read-intensive workloads, SELCC demonstrates great scalability regardless of the sharing ratio, as there is very little cache coherence overhead introduced in the system. For write-intensive and write-only workloads, SELCC scalability deteriorates with increased shared data ratio (Fig. 12c and d). The reason is that a higher shared data ratio increases the likelihood of two compute nodes caching the same data, resulting in a higher volume of invalidation messages. Compared to the fully partitioned SELCC (0% shared ratio), the fully shared SELCC (100% shared) shows a 29.5%/19.7% (8GB cache) performance degradation at 8 nodes in write-intensive and write-only workloads, respectively. However, SELCC still shows good scalability under write-intensive workloads. Compared to the single compute node deployment, the eight-node SELCC increases throughput by 6.24$\times$/5.42$\times$ corresponding to the write-intensive and write-only workloads.
@@ 7.1.2. Evaluating the Workloads with Access Locality
Workloads with access locality is one of the scenarios where the local cache can provide significant benefits. To illustrate the performance benefits of SELCC under workloads with access locality, we adjust the uniformly distributed workload so that each operation has a 50% probability of accessing the same GCL as the previous one. The benchmark is executed with 8 compute nodes, 100% sharing ratio, and varying numbers of threads across the nodes. Compared with SEL, due to the local cache, SELCC shows significant performance gains in all workloads (Fig. 13a), with improvements of 1.85$\times$/1.26$\times$/1.71$\times$ at 64 threads. Compared with GAM (TSO), GAM (SEQ) and ScaleStore, SELCC demonstrates higher performance in all four read ratios, achieving 5.57$\times$/6.31$\times$/1.67$\times$, 2.84$\times$/4.29$\times$/1.40$\times$, and 4.35$\times$/4.64$\times$/1.68$\times$ the throughput, respectively. GAM exhibits limited thread scalability due to its serialized queue for all the read/write operations and ScaleStore’s performance is limited by the insufficient remote computing power.
@@ 7.1.3. Evaluating the Workloads with Access Skewness
Workloads with access skewness represent another scenario where SELCC can achieve significant benefits. To illustrate the performance benefits of SELCC under a workload with access skewness, we run the benchmark with a Zipfian distribution, 100% sharing ratio. The skewness parameter, $\theta$, is set to 0.99, without applying the access locality. Other parameters are configured in the same way as that of the previous subsection. For read-intensive workloads, SELCC exhibits significant performance gains, achieving throughput 3.09$\times$ over that of SEL at 64 threads. These gains result from the high cache hit ratios (60.7%) of skewed workloads (Fig. 14). However, for write-intensive and write-only workloads, SEL shows better performance than SELCC initially when thread count is low, as SELCC suffers from a large number of invalidation messages triggered by the data hotspot. As the thread count increases, SEL experiences significant performance degradation (over 7$\times$ in write-intensive workloads), primarily due to the high contention in RDMA atomic operations over the data hotspot. In contrast, SELCC exhibits better thread scalability due to the prioritization of local concurrency control over global concurrency control (Sec.5.1.1), shifting the bottleneck from RDMA to local write-write conflict. Finally, SELCC outperforms GAM(TSO), GAM(SEQ) and ScaleStore by 4.74$\times$/5.85$\times$/2.09$\times$, 2.03$\times$/3.60$\times$/12.6$\times$, and 2.25$\times$/3.42$\times$/1.25$\times$ for workloads with 64 threads, highlighting the superiority of SELCC as a cache coherence protocol over disaggregated shared memory.
@@ 7.2. Evaluating Index Support over SELCC
Baselines. Four B-tree baselines are evaluated in this experiment. The first baseline is Sherman, an optimized index over disaggregated shared memory. The second baseline is the B-tree implementation over ScaleStore, configured the same as the one in the micro-benchmark. The third baseline is DEX, a sharding-based B-tree over disaggregated memory. Unlike the other shared-memory baselines, DEX employs a sharding mechanism to bypass the cache coherence problem. The final baseline is the B-tree over SEL.
Benchmarks & Configuration. We benchmark the indexes using YCSB, following methodologies established in the existing literature (Lu et al., 2024; Wang et al., 2023b, 2022). Each index is loaded with 2 billion key-value records (around 40GB) and tested under varying read ratios and data skewness ($\theta=0.99$). The experiments are conducted over 8 compute nodes, with 8 threads per node.
@@ 7.2.1. Results of Uniform Workloads
The B-tree over SELCC outperforms that over SEL by factors of 5.84$\times$/4.99$\times$/6.04$\times$/6.34$\times$, respectively (See Fig. 16a). Compared to Sherman, the B-tree over SELCC outperforms Sherman by factors of 1.63$\times$/1.42$\times$/1.69$\times$/1.76$\times$, respectively, because Sherman’s remote synchronization requires one more RDMA round trip compared with SELCC and Sherman cannot cache leaf nodes locally. SELCC outperforms scalestore by factor of 1.87$\times$/1.63$\times$/1.71$\times$/1.75$\times$, respectively, showing better performance as an abstraction layer over compute-limited disaggregated memory. Finally, the B-tree over SELCC slightly loses to DEX under uniform workloads. This result is expected, as the sharding mechanism in DEX fully bypasses the cache coherence problem and includes many index-specific optimizations. In contrast, the B-tree over SELCC serves as a demonstration with approximately 1,000 lines of code (LOC), significantly fewer than the typically extensive codebases of disaggregated indices, which often exceed 10,000 LOC. Moreover, it has limitations when serving as an index component in a full-fledged multi-primary database due to the overhead of cross-shard transactions (See Sec. 7.3).
@@ 8. Related Work
Database systems over disaggregated memory. Approaches to database research over disaggregated memory differ significantly between academia and industry. Academic research focuses on redesigning specific database components, e.g., indexes (Wang et al., 2022; Zuo et al., 2021; Luo et al., 2023; Wang et al., 2023b; Ziegler et al., 2019; Lu et al., 2024) and transaction concurrency control algorithms (Wei et al., 2018; Wang and Qian, 2021; Zamanian et al., 2016) over the disaggregated memory. SELCC converges the individual database research by providing a layer of abstraction. In contrast, industry, e.g., Alibaba PolarDB and Huawei GaussDB, conducts research in full-fledged system support over disaggregated memory (Cao et al., 2021; Zhang et al., 2021; Yang et al., 2024; Li et al., 2016; Ruan et al., 2023). They migrate the buffer pool onto disaggregated memory, achieving a higher cache hit ratio (Zhang et al., 2021; Cao et al., 2021), instant failure recovery (Cao et al., 2021; Li et al., 2016), elasticity resource provisioning (Cao et al., 2021), and multiple primary nodes (Yang et al., 2024; Li et al., 2024). Existing multi-primary databases use RPC-based protocols to maintain cache coherence over disaggregated memory, but their performance could be constrained by the limited remote computing power.

### v4

@@ 6.2. Consistency Model
By default, SELCC ensures sequential consistency. However, its consistency guarantees can be relaxed to enhance performance. For example, write operations can be buffered in a local queue and returned immediately, allowing execution order of reads and writes to be relaxed. These buffered write requests can then be processed in batches, enabling the multiplexing of invalidation messages and data transfers, which reduces RDMA bandwidth consumption. By carefully setting deadlines for processing the buffered write requests, SELCC can achieve varying levels of consistency, ranging from sequential consistency to eventual consistency.
@@ 6.3. Index Support
Leveraging the abstraction layer provided by SELCC, disaggregated indexes can be implemented with ease, either from scratch or by modifying existing codebases. For instance, we developed a concurrent B-link tree over SELCC using approximately 1,200 lines of code, of which around 200 lines related to SELCC, and modified an existing 1,400-line codebase to implement a concurrent R-tree with only 190 lines of changes. The primary steps in migrating an existing index to disaggregated memory involve aligning the index’s block with the Global Cache Line (GCL) and replacing mutex-based latches with SELCC latches. Nevertheless, not all index types can be migrated with such minimal modifications. In cases where the original data structure is not completely organized into blocks (e.g., bw-tree (Levandoski et al., 2013) and skip list (Pugh, 1990)), the data organization may require redesign; for instance, the sibling nodes in a skip list must be reorganized within the same GCL to avoid unnecessary RDMA operations during index traversal. Moreover, certain in-memory index optimizations, such as cache-conscious designs, remain effective in disaggregated environments, with the primary difference being that the CPU cache line size is replaced by the Global Cache Line size. We evaluate the concurrent B-link-tree’s performance using the YCSB benchmark (Fig. 17), finding it achieves competitive performance compared to state-of-the-art B-trees over shared disaggregated memory.
@@ 6.4. Transaction Support
We can migrate existing concurrency control algorithms to the disaggregated architecture by leveraging the abstraction layer enabled by SELCC as follows. (1) Tuples should be properly organized into GCLs. (2) Local shared-exclusive latches are replaced with SELCC_XLock/SELCC_SLock locks. (3) Algorithms that require monotonic timestamps utilize Atomic provided by SELCC API to perform RDMA Fetch-and-Add (FAA) operations on a global timestamp generator to obtain monotonically increasing timestamps.
We implement three types of algorithms over SELCC: Two-Phase Locking with no wait strategy (2PL), Timestamp Ordering (TO), and Optimistic Concurrency Control (OCC). Tuples are organized in a heap style, meaning they are placed in GCLs in chronological order of insertion. To ensure atomicity of tuple accesses, these accesses must be protected by SELCC_XLock/ SELCC_SLock locks. For two-phase locking, SELCC latches on the GCLs are reused for locking purposes, minimizing the RDMA round trips required by the transaction concurrency control. Since all transactions are executed within the same compute node via RDMA, transaction support over SELCC does not require two-phase commit protocols.
6.5. Discussion
To support a DBMS over disaggregated memory, current SELCC APIs exhibit several potential limitations, including data replication, failure recovery, and multi-tiered DB design. Despite various proposed solutions to data replication (Murat et al., 2024; Wang et al., 2023a; Behrens et al., 2018; Lee et al., 2022), the main challenge for SELCC memory is to incorporate these solutions without compromising performance. Failure recovery can leverage the two-tier ARIES protocol (Zhang et al., 2021), which addresses compute node and memory node failures differently. For memory node failures, data replication tolerates a limited number of faults; however, extensive node failures necessitate recovery using logs and checkpoints stored in disaggregated storage. In contrast, compute node recovery is accelerated by logs and checkpoints maintained in disaggregated memory, enabling surviving compute nodes to retrieve and replay logs based on these checkpoints. During failure recovery, dedicated threads scan the disaggregated memory to remove stale shared or exclusive ownerships associated with the failed compute nodes. Finally, when disaggregated memory serves as an extended buffer pool over disaggregated storage, maintaining global cache metadata (including the page address table and LRU list) imposes additional overhead on the critical path. A potential solution could involve managing these structures separately using an RPC-based protocol, thereby minimizing the extra RDMA round trips.
@@ 7.1. Evaluating SELCC
Baselines. To show the efficiency of SELCC, we compare SELCC against three competitors over disaggregated memory: (1) GAM (Cai et al., 2018), an RPC-based cache-coherence protocol designed for distributed shared memory. We test GAM with different consistency models: total store order consistency and sequential consistency, corresponding to GAM (TSO) and GAM (SEQ) in the figures. (2) ScaleStore (Ziegler et al., 2022), state-of-the-art RPC-based protocol designed for distributed shared memory over distributed shared SSDs. To adapt it for disaggregated memory, we configure it differently for compute and memory nodes. For the memory nodes, the buffer pool sizes are sufficiently large to fully store the data in memory, eliminating SSD access . For the compute nodes, the system no longer allocates data and instead functions as a stateless compute server. The last baseline is SEL (Ziegler et al., 2023b), a one-sided access framework that operates without compute-side caching. While it employs the shared-exclusive latch (SEL) to ensure RDMA access atomicity, it bypasses the cache coherence problem by disabling caching, which can lead to low performance.
@@
Benchmarks. We test the competitors by a micro-benchmark tool (Cai et al., 2018) that allows for adjustments in sharing ratios, read/write ratios, data skewness and access locality. The accesses in the micro-benchmark directly targets at the global address of GCL, removing the necessity of disaggregated index in the experiment. Each compute server issues 16 Million accesses over 24 Million allocated GCLs (2KB per GCL 4442KB GCL size reaches a good balancing between space and performance., 48GB in total). The overall throughput with different read ratios 100% (Read only), 95% (Read intensive), 50% (Write intensive), 0% (Write only) are tested. Read only workload has similar result to read intensive workload, so we omit its result in the presentation.
@@ 7.1.1. Evaluating the Scalability of SELCC
To evaluate the scalability of the SELCC protocol, we run the benchmark under a uniformly distributed workload while varying the number of compute nodes [^5] The total number of nodes is temporarily increased to 32 for this experiment.. We scale the number of memory nodes in proportion to the number of compute nodes. Furthermore, we increased the local cache size to 16GB to more clearly reveal the overhead associated with cache invalidation. We compare SELCC under various sharing ratios ($sr$), following the methodology in (Yang et al., 2024; Cai et al., 2018; Depoutovitch et al., 2023). The sharing ratio ($sr$) indicates the percentage of allocated data accessible by all compute nodes, while the remainder is accessed privately. When the sharing ratio is zero, the system essentially operates as a sharding-based system over disaggregated memory.
Fig. 12 shows the experimental results. The point values represent the overall throughput while the bar values indicate the proportion of operations requiring invalidation messages. For the read-intensive workloads, SELCC demonstrates great scalability regardless of the sharing ratio, as there is very little cache coherence overhead introduced in the system. For write-intensive and write-only workloads, SELCC scalability deteriorates with increased shared data ratio (Fig. 12c and d). The reason is that a higher shared data ratio increases the likelihood of two compute nodes caching the same data, resulting in a higher volume of invalidation messages. Compared to the fully partitioned SELCC (0% shared ratio), the fully shared SELCC (100% shared) shows a 37.7%/33.2% (8GB cache) performance degradation at 16 nodes in write-intensive and write-only workloads, respectively. However, SELCC still shows good scalability under write-intensive workloads. Compared to the single compute node deployment, the 16-node SELCC increases throughput by 10.4$\times$/9.83$\times$ corresponding to the write-intensive and write-only workloads.
@@ 7.1.2. Evaluating the Workloads with Access Locality
Workloads with access locality is one of the scenarios where the local cache can provide significant benefits. To illustrate the performance benefits of SELCC under workloads with access locality, we adjust the uniformly distributed workload so that each operation has a 50% probability of accessing the same GCL as the previous one. The benchmark is executed with 8 compute nodes, 100% sharing ratio, and varying numbers of threads across the nodes. Compared with SEL, due to the local cache, SELCC shows significant performance gains in all workloads (Fig. 13a), with improvements of 1.85$\times$/1.26$\times$/1.71$\times$ at 64 threads. Compared with GAM (TSO), GAM (SEQ) and ScaleStore, SELCC demonstrates higher performance in all four read ratios, achieving 5.57$\times$/6.31$\times$/1.67$\times$, 2.84$\times$/4.29$\times$/1.40$\times$, and 4.35$\times$/4.64$\times$/1.68$\times$ the throughput, respectively. GAM exhibits limited thread scalability due to its serialized queue for all the read/write operations and ScaleStore’s performance is limited by the insufficient remote computing power. Additionally, we executed the pure-write workload on SELCC with varying GCL sizes and observed that throughput decreases as the GCL size increases (See Table below). This behavior is attributable to the fact that larger GCL sizes result in increased RDMA payloads, which in turn prolong the latency of RDMA operations.
512B 1KB 2KB 4KB 8KB
Throughput (Mops/s) 8.82 7.70 7.11 5.81 4.30
@@ 7.1.3. Evaluating the Workloads with Access Skewness
Workloads with access skewness represent another scenario where SELCC can achieve significant benefits. Moreover, workload with extreme access skewness can trigger unfairness or even starvation of data access.
To illustrate the performance benefits of SELCC under a workload with access skewness, we run the benchmark with a Zipfian distribution, 100% sharing ratio. The skewness parameter, $\theta$, is set to 0.99, without applying the access locality. Other parameters are configured in the same way as that of the previous subsection. For read-intensive workloads, SELCC exhibits significant performance gains, achieving throughput 3.09$\times$ over that of SEL at 64 threads. These gains result from the high cache hit ratios (60.7%) of skewed workloads (Fig. 14). However, for write-intensive and write-only workloads, SEL shows better performance than SELCC initially when thread count is low, as SELCC suffers from a large number of invalidation messages triggered by the data hotspot. As the thread count increases, SEL experiences significant performance degradation (over 7$\times$ in write-intensive workloads), primarily due to the high contention in RDMA atomic operations over the data hotspot. In contrast, SELCC exhibits better thread scalability due to the prioritization of local concurrency control over global concurrency control (Sec.5.1.1), shifting the bottleneck from RDMA to local write-write conflict. Finally, SELCC outperforms GAM(TSO), GAM(SEQ) and ScaleStore by 4.74$\times$/5.85$\times$/2.09$\times$, 2.03$\times$/3.60$\times$/12.6$\times$, and 2.25$\times$/3.42$\times$/1.25$\times$ for workloads with 64 threads, highlighting the superiority of SELCC as a cache coherence protocol over disaggregated shared memory.
In order to illustrate the effective of our approahches in improving fairness, we employ two distinct experimental setups. First, we run a write-only workload with varying handover thresholds $\gamma$ as introduced in Sec. 5.3.1. Second, we configure a single-writer, multiple-readers setup, in which one compute node executes a write-only workload while the others execute read-only workloads. An ablation study of the spin-waiting and priority-matching mechanisms (detailed in Section 5.3.2) is conducted in the setup. In both experiments, we set the skewness parameter $\theta$ to 10. As shown in Fig. 15 a, prioritizing local ownership transfer ($\gamma=Inf$) over global ownership transfer ($\gamma=0$) improves performance but reduces access fairness. Conversely, selecting an intermediate threshold ($\gamma=256$) effectively balances performance and fairness  [^6] Under highly skewed pure-write workload, the amortized latency per data access is $\frac{CT_{rm}}{\gamma}+CPT_{lm}$, where $C$ denotes the number of compute nodes, P is the thread number per compute node, $T_{rm}$ represents the latency of an RDMA roundtrip, $T_{lm}$ is the latency of main memory access. . For the single-writer, multiple-readers experiment, we depict the cumulative executed operations over time for one reader node and one writer node. When no optimization is applied, the writer experiences starvation (Fig. 15 b). After enabling spin-waiting, the writer is no longer blocked by the readers, but its throughput remains substantially lower. Once we activate the priority-matching mechanism, the writer’s performance becomes comparable to that of the readers, demonstrating significantly enhanced access fairness.
@@
Refer to caption Refer to caption
(a) Performance per node under write-only workload. (b) Cumulative operations over time for single-writer and multiple-readers setup.
Figure 15. Performance evaluation on access fairness.
@@ 7.2. Evaluating Index Support over SELCC
Baselines. Four B-tree baselines are evaluated in this experiment. The first baseline is Sherman (Wang et al., 2022), an optimized index over disaggregated shared memory. The second baseline is the B-tree implementation over ScaleStore (Ziegler et al., 2022), configured the same as the one in the micro-benchmark. The third baseline is DEX (Lu et al., 2024), a sharding-based B-tree over disaggregated memory. Unlike the other shared-memory baselines, DEX employs a sharding mechanism to bypass the cache coherence problem. The final baseline is the B-tree over SEL.
Benchmarks & Configuration. We benchmark the indexes using YCSB, following methodologies established in the existing literature (Lu et al., 2024; Wang et al., 2023c, 2022). Each index is loaded with 2 billion key-value records (around 40GB) and tested under varying read ratios and data skewness ($\theta=0.99$). The experiments are conducted over 8 compute nodes, with 8 threads per node.
@@ 7.2.1. Results of Uniform Workloads
The B-tree over SELCC outperforms that over SEL by factors of 5.84$\times$/4.99$\times$/6.04$\times$/6.34$\times$, respectively (See Fig. 17a). Compared to Sherman, the B-tree over SELCC outperforms Sherman by factors of 1.63$\times$/1.42$\times$/1.69$\times$/1.76$\times$, respectively, because Sherman’s remote synchronization requires one more RDMA round trip compared with SELCC and Sherman cannot cache leaf nodes locally. SELCC outperforms scalestore by factor of 1.87$\times$/1.63$\times$/1.71$\times$/1.75$\times$, respectively, showing better performance as an abstraction layer over compute-limited disaggregated memory. Finally, the B-tree over SELCC slightly loses to DEX under uniform workloads. This result is expected, as the sharding mechanism in DEX fully bypasses the cache coherence problem and includes many index-specific optimizations. In contrast, the B-tree over SELCC serves as a demonstration with approximately 1,000 lines of code (LOC) over SELCC, significantly fewer than the typically extensive codebases of disaggregated indices (Wang et al., 2022; Zuo et al., 2021; Ziegler et al., 2019; Wang et al., 2023c; Luo et al., 2023). Moreover, DEX has limitations when serving as an index component in a full-fledged multi-primary database due to the overhead of cross-shard transactions (See Sec. 7.3).
@@ 8. Related Work
Database systems over disaggregated memory. Approaches to database research over disaggregated memory differ significantly between academia and industry. Academic research focuses on redesigning specific database components, e.g., indexes (Wang et al., 2022; Zuo et al., 2021; Luo et al., 2023; Wang et al., 2023c; Ziegler et al., 2019; Lu et al., 2024) and transaction concurrency control algorithms (Wei et al., 2018; Wang and Qian, 2021; Zamanian et al., 2016) over the disaggregated memory. SELCC converges the individual database research by providing a layer of abstraction. In contrast, industry, e.g., Alibaba PolarDB and Huawei GaussDB, conducts research in full-fledged system support over disaggregated memory (Cao et al., 2021; Zhang et al., 2021; Yang et al., 2024; Li et al., 2016; Ruan et al., 2023). They migrate the buffer pool onto disaggregated memory, achieving a higher cache hit ratio (Zhang et al., 2021; Cao et al., 2021), instant failure recovery (Cao et al., 2021; Li et al., 2016), elasticity resource provisioning (Cao et al., 2021), and multiple primary nodes (Yang et al., 2024; Li et al., 2024). Existing multi-primary databases use RPC-based protocols to maintain cache coherence over disaggregated memory, but their performance could be constrained by the limited remote computing power.